If h(x) = 6, for the function h(x) = 2x + 1, what is the value of x
Found 2 solutions by Cromlix, fractalier:
Answer by Cromlix(4381)

Hi there,
If h(x) = 6, for the function h(x) = 2x + 1
2x + 1 = 6
2x = 6 - 1
2x = 5
x = 5/2 or 2.5
Hope this helps :-)

Answer by fractalier(6550)

For h(x) = 2x + 1 = 6, we merely solve for x...
2x = 5
x = 5/2
such that
h(5/2) = 6